I am stuck on correct answer for:

if a triangle had these degrees of measurement what would be the degree measure for B?
A = left side of trangle is degree measure of 30
C = top of triangle is degree measure of 100
B = ?

What is correct resolve to this?

Answer by venugopalramana(3286)

I am stuck on correct answer for:
if a triangle had these degrees of measurement what would be the degree measure for B?
A = left side of trangle is degree measure of 30
C = top of triangle is degree measure of 100
B = ?
What is correct resolve to this?
SUM OF 3 ANGLES IN A TRIANGLE = 180
HERE SUM OF 2 ANGLES = 30+100=130
HENCE THE THIRD ANGLE = B = 180-130 = 50